Sir please describe me physical meaning of angular momentum

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Consider a particle capable of rotation about an axis. At any time t, let its momentum be and , the position vector. Then the angular momentum is given by the cross product and . = x .

In three dimensions,

Lx = ypz - zpy, Ly = zpx - xpz, Lz = xpy - ypx.

The angular momentum, which is nothing but moment of linear momentum, can be expressed in terms of the lever arm for momentum.

Let us consider the case of angular momentum in two dimensions.

L = xpy - ypx

If  j is the angle between and x-axis, then px = p cos y and py = p sin y. Also x = r cos q, y = r sin q.

Hence, we can define angular momentum as the product of linear momentum and the lever arm for momentum.

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi